OpenStudy (anonymous):

Please help :/ Express 2sin3theta sin5theta as an algebraic sum.

OpenStudy (anonymous):

Use: \[\large \color{green}{2\sin(\frac{C+D}{2})\sin(\frac{D-C}{2}) = \cos(C) - \cos(D)}\]

OpenStudy (anonymous):

Here: \[\frac{C+D}{2} = 3\] \[\frac{D-C}{2} = 5\] Solve them to find C and D here..

OpenStudy (anonymous):

Ok, just a sec

OpenStudy (anonymous):

\[C + D = 6\] \[-C + D =1 0\] Add them now you will get D..

OpenStudy (anonymous):

16

OpenStudy (anonymous):

Left hand side it will be: 2D = 16 Now find D from here..

OpenStudy (anonymous):

8?

OpenStudy (anonymous):

Yes.. Now you have D so find C now by plugging in D value in any one equation...

OpenStudy (anonymous):

its looks like that double angle formula can be used.

OpenStudy (anonymous):

\[C + D = 16 \implies C = 6 - D \implies C = 6 - 8 \implies C = -2\]

OpenStudy (anonymous):

@muhammad9t5 which double angle will you use here ??

OpenStudy (anonymous):

And do you know what question wants?/ It wants Algebraic Sum.. Meaning Plus or Minus separated terms..

OpenStudy (anonymous):

So you will have now C = -2 and D = 8 \[\large \cos(-2) - \cos(8) \implies \color{green}{\cos(2) - \cos(8)}\] \[\cos(- \theta) = \cos(\theta)\]

OpenStudy (anonymous):

This will help. Here are the options: a. -cos8theta+cos2theta b. cos8theta-cos2theta c. -cos2theta-cos2theta d. cos8theta+cos2theta

OpenStudy (anonymous):

How about first one ??

OpenStudy (anonymous):

Yes, that's what I was thinking. @muhammad9t5 can you confirm?

OpenStudy (anonymous):

\[\large \cos(-2\theta) - \cos(8 \theta) \implies \color{green}{\cos(2\theta) - \cos(8\theta)}\]

OpenStudy (anonymous):

I confirm !

OpenStudy (anonymous):

Yes, @waterineyes, perfect. hanks @Neemo. Thanks to everyyone who helped!

OpenStudy (anonymous):

Welcome dear.. Interact when you are asked something @IloveCharlie

OpenStudy (anonymous):

yw !

OpenStudy (anonymous):

@waterineyes sorry it was sum to product identity..